What is the future value of 25 periodic payments of $5,630 each made at the beginning of each period and compounded at 8%?

What is the present value of $3,440 to be received at the beginning of each of 29 periods, discounted at 5% compound interest?

What is the future value of 15 deposits of $2,640 each made at the beginning of each period and compounded at 10%? (Future value as of the end of the 15th period.)

What is the present value of 7 receipts of $3,140 each received at the beginning of each period, discounted at 9% compounded interest?

What is the future value of $7,890 at the end of 5 periods at 8% compounded interest?

What is the present value of $7,890 due 9 periods hence, discounted at 6%?

What is the future value of 16 periodic payments of $7,890 each made at the end of each period and compounded at 10%?

What is the present value of $7,890 to be received at the end of each of 19 periods, discounted at 5% compound interest?
Answer #1

1.
=5630/8%*(1.08^25-1)*1.08
=444513.3573

2.
=3440/5%*(1-1/1.05^29)*1.05
=54689.55776

3.
=2640/10%*(1.1^15-1)*1.1
=92267.28684

4.
=3140/9%*(1-1/1.09^7)*1.09
=17225.78437